# A Differentiable Model for Optimizing the Genetic Drivers of Synaptogenesis

### v2

@@ 4 Discussion
Moreover, the current implementation of SynaptoGen assumes a feedforward multipartite network topology, which can be achieved by genetically inhibiting, through an additional set of genes, all synapses that could form between compatible neurons in non-adjacent layers. Another possiblity would be to employ an external pruning process. However, biological neuronal networks exhibit more complex, often cyclic wiring. SynaptoGen can be extended to account for this complexity by integrating its weight matrix decomposition into frameworks like 4Ward (Boccato et al., 2024a; b), which are designed to convert arbitrary graphs into neural networks trainable via backpropagation. Incorporating more complex topologies would also enable compliance with constraints derived from contactomic data (i.e., information about the physical contacts between neurons), as synapses can only form between neurons in close proximity. Using such topologies, SynaptoGen could learn the average number of synapses only for neuron pairs which are actually in contact. Another phenomenon which his not modeled in the current version of SynaptoGen is the modulation of genetic rules based on the distance between neurons. As highlighted in (Stöckl et al., 2022), connection probabilities decay exponentially with increasing distance. This behavior can be incorporated into SynaptoGen without compromising differentiability by assigning each neuron a positional embedding to calculate distances. This embedding could be learnable, given as a prior, or derived from genetic expression. We recall that a function such as $e^{-||\bm{p}^{u}-\bm{p}^{v}||}$, where $\bm{p}^{u}$ and $\bm{p}^{v}$ are the positions of neurons $u$ and $v$, is differentiable w.r.t. the positions except at coincident points.
@@ Appendix A The Polarity Matrix
Here, each neurotransmitter synthesized in pre-synaptic neurons can be bind to either a positive (+) or negative (-) receptor in post-synaptic neurons. The 0s in the matrix signify that receptors attuned to a specific neurotransmitter are incapable of receiving different ones. This formalism readily extends to accommodate an arbitrary number of neurotransmitters by setting $M=2L$ and expanding $A$ to an $L\times 2L$ block diagonal matrix, where each block is represented as $[1,-1]$. It is worth noting that, in the experiments described in Section 3, the entries of $A$ do not belong to the set of learnable parameters.
@@ Appendix B Simulating Synaptogenesis
Algorithm 1 outlines one possible implementation of the synaptogenesis simulation introduced at the end of Section 2, which we used in our experiments.
@@
The coefficient $\alpha$ is a general correction factor used to control the average degree of nodes in the generated networks. The necessity of this correction is discussed in Appendix C. Here, we show that this correction does not impact the average weight matrices learned by SynaptoGen:
@@
In all our simulations, we chose $\alpha$ such that the resulting networks had an average degree of $10^{5}$. While this value is quite high (the typical number of synapses per neuron in the human brain is on the order of magnitude of $10^{4}$ (Zhang, 2019)), it aligns with the connectivity levels observed in certain types of cells in specific brain regions, e.g., Purkinje cells in the cerebellar cortex (Napper & Harvey, 1988).
@@ Appendix C A Problem of Quantization
In this proof, $n_{ij}p_{ij}=\mathbb{E}[\text{Bin}(n_{ij},p_{ij})]$, the outer round in step equation 15 helps account for the nearest integer approximation of the mean number of synapses learned by SynaptoGen, and steps equation 17 and equation 18 are derived from bounds applied to the two rounding operations.
@@ Appendix D Training Details
This section provides all the details related to the training of the SynaptoGen models that were not explicitly mentioned in Section 2 Neural networks were implemented as custom MLPs featuring a single hidden layer with 128 neurons, a number that allowed us to easily conduct experiments with the computational resources available. The input and output layers were configured with a number of neurons equal to the dimensionality of the observation space and the number of possible actions in the selected RL environments, respectively. Consequently, the size of the input layer ranged from 2 to 8, while the size of the output layer ranged from 2 to 4. Details about the RL environments can be found in the Gym library documentation: https://www.gymlibrary.dev/index.html. Learnable parameters associated with genetic rules and conductances were initialized following the procedure described in (Barabási et al., 2023). All other parameters were initialized using a Kaiming normal distribution (He et al., 2015).
Trainings were conducted using the Adam optimizer (Kingma & Ba, 2015), with its default parameters unchanged except for the learning rate. Training consisted of 500000 environment steps. To minimize randomness in the results and mitigate sub-optimality associated with specific hyperparameter choices, the genetic profiles used in simulations were selected through a hyperparameter grid-search. This search explored three different learning rates (0.03, 0.003, 0.0003) and three random seeds for various sampling processes. Specifically, after hyperparameter optimization, we retained the checkpoint corresponding to the best validation performance achieved by the models. Validation was conducted every 10000 training steps. During both validation and subsequent simulations, agent rewards were computed as the average scores over 10 different episodes. The remaining hyperparameters related to the DQN algorithm were sourced from: https://github.com/DLR-RM/rl-baselines3-zoo/blob/master/hyperparams/dqn.yml. Additional information can be found in the released code.
@@ Appendix E The Separable Natural Evolution Strategy (SNES) Baseline
As mentioned in Section 3, the SNES baseline was constructed by replacing gradient descent in our SynaptoGen framework with SNES, an optimization technique introduced in (Schaul, 2012) and employed by the authors of (Stöckl et al., 2022) for a related task, namely the optimization of probabilistic skeletons. The core idea behind SNES is to sample $\lambda$ offsprings from a normal distribution of the type $\mathcal{N}(\bm{\mu},I\bm{\sigma})$, where $\bm{\mu}$ is built by concatenating the flattened learnable matrices of SynaptoGen. These offsprings are evaluated based on their fitness, allowing the distribution to adapt and better capture those regions of the parameter space where offspring fitness is higher. The implementation details of SNES are provided in the pseudo-code of Algorithm 2.
@@
while num_steps $<$ max_steps do
@@
       num_steps += $steps\_performed()$
@@
Inside the algorithm, the vectors $\bm{s}_{k}$ contain the “noise” used in generating the offsprings, $\bm{\sigma}$ is a vector of variances initialized in accordance with (Stöckl et al., 2022), $F$ is the fitness function, $\eta_{\bm{\mu}}$ and $\eta_{\bm{\sigma}}$ are two learning rates $>0$, and the method $steps\_performed()$ returns the number of actions taken during the current iteration on the RL environment being optimized. Specifically, we defined the fitness function $F(\bullet)$ as the average performance achieved by $m$ simulated biological agents based on the genetic profile provided as input to the function. Individual scores were calculated by averaging the rewards obtained by an agent over 10 different episodes.
To ensure fair comparisons with the profiles obtained through gradient descent, we limited the number of actions executable on a RL environment to 500000 in the context of a single optimization. Additionally, for this baseline, we conducted a hyperparameter search similar to the one described in Appendix D, varying $m$ in the set $\{10,20,30\}$ and assigning $\lambda$, during the search, the values 8, 16, and the default value of the SNES implementation used (https://github.com/pybrain/pybrain).
@@ Appendix F The Bio-Plausible Baseline
In more detail, for each agent, we sampled the genetic expression profile of a virtual zygote from a normal distribution and simulated a lineage tree with as many leaves as the number of neurons in the agent. This allowed us to assign gene expression profiles to individual cells following the equations above. We finalized the procedure by randomly assigning each neural layer expression profiles corresponding to leaves from the same complete subtree, simulating spatial contiguity for neurons within the same layer. The procedure we implemented is detailed in Algorithm 3. In the pseudo-code, the first line initializes the zygote’s genetic expression, $N$ corresponds to the target number of cells to be produced, the notation $\bullet_{1\dots N}$ selects the first $N$ columns of the respective matrix, function $rand\_roll()$ circularly permutes the columns of the input matrix, and $C^{in}$, $C^{hidden}$ and $C^{out}$ represent the matrices associated with the genetic expressions of the various layers of the agent being initialized.
@@
       num_cells $\leftarrow$ $2$num_cells
@@ Appendix H Additional Results
To produce the genetic rules used in the bio-plausible validation in Section 3, we first computed the gene co-expressions of the C. elegans nerve ring by performing a generalized nDGE analysis on the dataset provided in (Taylor et al., 2021). This generalization was necessary because the nDGE implementation available at https://github.com/cengenproject/connectivity_analysis is designed to calculate local co-expressions, specifically limited to a set of cells consisting of a pre-synaptic neuron and its corresponding post-synaptic partners.
To compute global co-expressions, we identified all pairs of neurons that shared synapses and all the pairs that were in contact without being connected. Then, for each possible pair of genes, we calculated co-expressions for both sets and conducted a T-test with Bonferroni correction. Pairs with a $p$-value $<0.05$ were considered co-expressed. The detailed procedure is outlined in Algorithm 4. In this pseudo-code, $B$ and $C$ represent the connectome and the contactome of the neuronal network under analysis, respectively, $\oplus$ denotes a pointwise XOR operation, and $G$ corresponds to the number of genes analyzed. The third-level “for” loops, instead, separate the indices of pre- and post-synaptic neurons involved in the pairs being evaluated in each iteration. The co-expression of a specific pair of genes for all neuron pairs in a given set is calculated as the pointwise product $C_{i,\bm{u}}\odot C_{j,\bm{v}}$, where $C$ is the matrix containing gene expression data for all neurons, and the notation $\bullet_{i,\bm{u}}$ selects the $i$-th row (corresponding to gene $i$) and the columns associated with the neurons in $\bm{u}$. The method $ttest()$ performs the T-test, while $p_{ij}$ and $g_{ij}$ represent the $p$-value for the gene pair $(i,j)$ and its corresponding binary co-expression, respectively.
@@
             $p_{ij}\leftarrow ttest($coexp_ws, coexp_co))))

### v3

@@ 4 Discussion
[…]
@@ 4.1 Conclusions
5 Acknowledgments
This work is supported and funded by: #NEXTGENERATIONEU (NGEU); the Ministry of University and Research (MUR); the National Recovery and Resilience Plan (NRRP); project MNESYS (PE0000006, to NT) - A Multiscale integrated approach to the study of the nervous system in health and disease (DN. 1553 11.10.2022); the MUR-PNRR M4C2I1.3 PE6 project PE00000019 Heal Italia (to NT); the NATIONAL CENTRE FOR HPC, BIG DATA AND QUANTUM COMPUTING, within the spoke “Multiscale Modeling and Engineering Applications” (to NT); the European Innovation Council (Project CROSSBRAIN - Grant Agreement 101070908, Project BRAINSTORM - Grant Agreement 101099355); the Horizon 2020 research and innovation Programme (Project EXPERIENCE - Grant Agreement 101017727). Tommaso Boccato is a PhD student enrolled in the National PhD in Artificial Intelligence, XXXVII cycle, course on Health and Life Sciences, organized by Università Campus Bio-Medico di Roma.
@@ Appendix A The Polarity Matrix
Here, each neurotransmitter synthesized in pre-synaptic neurons can be bind to either a positive (+) or negative (-) receptor in post-synaptic neurons. The 0s in the matrix signify that receptors attuned to a specific neurotransmitter are incapable of receiving different ones. This formalism readily extends to accommodate an arbitrary number of neurotransmitters by setting $M=2L$ and expanding $A$ to an $L\times 2L$ block diagonal matrix, where each block is represented as [1,1][1,-1]. It is worth noting that, in the experiments described in Section Results, the entries of $A$ do not belong to the set of learnable parameters.
@@ Appendix B Simulating Synaptogenesis
Algorithm 1 outlines one possible implementation of the synaptogenesis simulation introduced at the end of Section Methods, which we used in our experiments.
@@
The coefficient $\alpha$ is a general correction factor used to control the average degree of nodes in the generated networks. The necessity of this correction is discussed in Appendix A Problem of Quantization. Here, we show that this correction does not impact the average weight matrices learned by SynaptoGen:
@@
In all our simulations, we chose $\alpha$ such that the resulting networks had an average degree of $10^{5}$. While this value is quite high (the typical number of synapses per neuron in the human brain is on the order of magnitude of $10^{4}$ (Zhang, 2019)), it aligns with the connectivity levels observed in certain types of cells in specific brain regions, e.g., Purkinje cells in the cerebellar cortex (Napper and Harvey, 1988).
@@ Appendix C A Problem of Quantization
In this proof, $n_{ij}p_{ij}=\mathbb{E}[\text{Bin}(n_{ij},p_{ij})]$, the outer round in step (15) helps account for the nearest integer approximation of the mean number of synapses learned by SynaptoGen, and steps (17) and (18) are derived from bounds applied to the two rounding operations.
@@ Appendix D Training Details
This section provides all the details related to the training of the SynaptoGen models that were not explicitly mentioned in Section Methods Neural networks were implemented as custom MLPs featuring a single hidden layer with 128 neurons, a number that allowed us to easily conduct experiments with the computational resources available. The input and output layers were configured with a number of neurons equal to the dimensionality of the observation space and the number of possible actions in the selected RL environments, respectively. Consequently, the size of the input layer ranged from 2 to 8, while the size of the output layer ranged from 2 to 4. Details about the RL environments can be found in the Gym library documentation: https://www.gymlibrary.dev/index.html. Learnable parameters associated with genetic rules and conductances were initialized following the procedure described in (Barabási et al., 2023). All other parameters were initialized using a Kaiming normal distribution (He et al., 2015).
Trainings were conducted using the Adam optimizer (Kingma and Ba, 2015), with its default parameters unchanged except for the learning rate. Training consisted of 500000 environment steps. To minimize randomness in the results and mitigate sub-optimality associated with specific hyperparameter choices, the genetic profiles used in simulations were selected through a hyperparameter grid-search. This search explored three different learning rates (0.03, 0.003, 0.0003) and three random seeds for various sampling processes. Specifically, after hyperparameter optimization, we retained the checkpoint corresponding to the best validation performance achieved by the models. Validation was conducted every 10000 training steps. During both validation and subsequent simulations, agent rewards were computed as the average scores over 10 different episodes. The remaining hyperparameters related to the DQN algorithm were sourced from: https://github.com/DLR-RM/rl-baselines3-zoo/blob/master/hyperparams/dqn.yml. Additional information can be found in the released code.
@@ Appendix E The Separable Natural Evolution Strategy (SNES) Baseline
As mentioned in Section Results, the SNES baseline was constructed by replacing gradient descent in our SynaptoGen framework with SNES, an optimization technique introduced in (Schaul, 2012) and employed by the authors of (Stöckl et al., 2022) for a related task, namely the optimization of probabilistic skeletons. The core idea behind SNES is to sample $\lambda$ offsprings from a normal distribution of the type $\mathcal{N}(\bm{\mu},I\bm{\sigma})$, where $\bm{\mu}$ is built by concatenating the flattened learnable matrices of SynaptoGen. These offsprings are evaluated based on their fitness, allowing the distribution to adapt and better capture those regions of the parameter space where offspring fitness is higher. The implementation details of SNES are provided in the pseudo-code of Algorithm 2.
@@
while num_steps << max_steps do
@@
   num_steps += steps_performed()steps\_performed()
@@
Inside the algorithm, the vectors $\bm{s}_{k}$ contain the “noise” used in generating the offsprings, $\bm{\sigma}$ is a vector of variances initialized in accordance with (Stöckl et al., 2022), $F$ is the fitness function, $\eta_{\bm{\mu}}$ and $\eta_{\bm{\sigma}}$ are two learning rates >0>0, and the method steps_performed()steps\_performed() returns the number of actions taken during the current iteration on the RL environment being optimized. Specifically, we defined the fitness function $F(\bullet)$ as the average performance achieved by $m$ simulated biological agents based on the genetic profile provided as input to the function. Individual scores were calculated by averaging the rewards obtained by an agent over 10 different episodes.
To ensure fair comparisons with the profiles obtained through gradient descent, we limited the number of actions executable on a RL environment to 500000 in the context of a single optimization. Additionally, for this baseline, we conducted a hyperparameter search similar to the one described in Appendix Training Details, varying $m$ in the set $\{10,20,30\}$ and assigning $\lambda$, during the search, the values 8, 16, and the default value of the SNES implementation used (https://github.com/pybrain/pybrain).
@@ Appendix F The Bio-Plausible Baseline
In more detail, for each agent, we sampled the genetic expression profile of a virtual zygote from a normal distribution and simulated a lineage tree with as many leaves as the number of neurons in the agent. This allowed us to assign gene expression profiles to individual cells following the equations above. We finalized the procedure by randomly assigning each neural layer expression profiles corresponding to leaves from the same complete subtree, simulating spatial contiguity for neurons within the same layer. The procedure we implemented is detailed in Algorithm 3. In the pseudo-code, the first line initializes the zygote’s genetic expression, $N$ corresponds to the target number of cells to be produced, the notation $\bullet_{1\dots N}$ selects the first $N$ columns of the respective matrix, function rand_roll()rand\_roll() circularly permutes the columns of the input matrix, and $C^{in}$, $C^{hidden}$ and $C^{out}$ represent the matrices associated with the genetic expressions of the various layers of the agent being initialized.
@@
   num_cells $\leftarrow$ 22num_cells
@@ Appendix H Additional Results
To produce the genetic rules used in the bio-plausible validation in Section Results, we first computed the gene co-expressions of the C. elegans nerve ring by performing a generalized nDGE analysis on the dataset provided in (Taylor et al., 2021). This generalization was necessary because the nDGE implementation available at https://github.com/cengenproject/connectivity_analysis is designed to calculate local co-expressions, specifically limited to a set of cells consisting of a pre-synaptic neuron and its corresponding post-synaptic partners.
To compute global co-expressions, we identified all pairs of neurons that shared synapses and all the pairs that were in contact without being connected. Then, for each possible pair of genes, we calculated co-expressions for both sets and conducted a T-test with Bonferroni correction. Pairs with a $p$-value <0.05<0.05 were considered co-expressed. The detailed procedure is outlined in Algorithm 4. In this pseudo-code, $B$ and $C$ represent the connectome and the contactome of the neuronal network under analysis, respectively, $\oplus$ denotes a pointwise XOR operation, and $G$ corresponds to the number of genes analyzed. The third-level “for” loops, instead, separate the indices of pre- and post-synaptic neurons involved in the pairs being evaluated in each iteration. The co-expression of a specific pair of genes for all neuron pairs in a given set is calculated as the pointwise product $C_{i,\bm{u}}\odot C_{j,\bm{v}}$, where $C$ is the matrix containing gene expression data for all neurons, and the notation $\bullet_{i,\bm{u}}$ selects the $i$-th row (corresponding to gene $i$) and the columns associated with the neurons in $\bm{u}$. The method $ttest()$ performs the T-test, while $p_{ij}$ and $g_{ij}$ represent the $p$-value for the gene pair $(i,j)$ and its corresponding binary co-expression, respectively.
@@
    $p_{ij}\leftarrow ttest($coexp_ws, coexp_co))